Size of government question. Find Statistics Canada Table: 14-10-0027- 01. Include that table and the calculations on the sheet named Prob 2 in the template.
a. From that Table, establish the ratio of public sector employees to private sector employees (do not include the self-employed category) in 1987 and then compare this ratio with that of 2019. How do they compare?

b. Public sector employment typically consists of three groups: 1) Educational services; 2) Health care and social assistance; and Public administration. By how much in % has each one of these groups increased from 1987 to 2019?

c. Compare private sector employment with public sector employment. Which one has increased in % the most over the 1987 to 2019 period?
Answer #1

1.

ratio of public sector employees to private sector employees (1987)=2625/8008.9=.3278

ratio of public sector employees to private sector employees (2019)=3,849.3/12,303.7=.3129

The Ratio has decreased, meaning higher percentage of people are employed in private sector now.

2.

Count (1987,'000) Count (2019,'000) Growth
Educational Services 698.7 1160.2 66.05%
Health care and social assistance 635 1110.5 74.88%
Public Administration 756.7 1011.3 33.65%

3.

Count (1987,'000) Count (2019,'000) Growth
Public Sector 2625 3849.3 46.64%
Private Sector 8008.9 12303.7 53.63%

Private sector increased the most.